# WaveFormer: A Lightweight Transformer Model for sEMG-based Gesture Recognition

### v2

@@ V CONCLUSIONS
The sEMG datasets used in this study (EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG) were collected following approved ethical protocols. The original data collection procedures involving human subjects were approved by the respective Institutional Review Boards of the institutions where the data were collected. All participants provided informed consent, and the experimental procedures followed the principles outlined in the Declaration of Helsinki. No additional human subject experiments were conducted specifically for this study.
@@ ACKNOWLEDGMENT
This research was partly supported by the EU Horizon Europe project IntelliMan (g.a. 101070136), the PNRR MUR project ECS00000033ECOSISTER, and the ETH Zürich’s Future Computing Laboratory funded by a donation from Huawei Technologies. Furthermore, the research was also partially supported by the EU Horizon Europe project HAL4SDV (g.a. 101139789).

### v3

@@ V CONCLUSIONS
The sEMG datasets used in this study were collected following approved ethical protocols. Original data collection procedures involving human subjects were approved by the respective Institutional Review Boards of the institutions where the data was collected. All participants provided their informed consent and the experimental procedures followed the principles outlined in the Declaration of Helsinki. No additional human subject experiments were performed specifically for this study.
@@ ACKNOWLEDGMENT
This research was supported by the EU Horizon Europe project IntelliMan (grant agreement 101070136), the PNRR MUR project ECS00000033 ECOSISTER, the ETH Zürich’s Future Computing Laboratory funded by a donation from Huawei Technologies, and also the EU Horizon Europe project HAL4SDV (grant agreement 101139789). In addition, we gratefully acknowledge the computational resources made available by Cineca and CSCS. In particular, this work utilised the GPU-based high-performance computing infrastructures provided by Cineca and CSCS, whose support and operational services were instrumental in enabling the large-scale simulations/data-analysis undertaken for this study.